OpenStudy (anonymous):

Integration help! integral of sec^5 x dx

OpenStudy (anonymous):

/barf im guessing u sub will help you here

OpenStudy (across):

Integration by parts!

OpenStudy (across):

Whoops, I misread.

OpenStudy (anonymous):

yes integration by parts

OpenStudy (anonymous):

(sec^2x)(sec^2)(sec x)

OpenStudy (across):

I read it as sec^5(x)x for a moment there. Hmm, in this case, you have to use a reduction formula.

OpenStudy (anonymous):

red formula will work if you have it

OpenStudy (anonymous):

well i havent learned taht so is tehre anoterh way to do it by parts

OpenStudy (anonymous):

sec^2x = 1 + tan ^2

OpenStudy (anonymous):

( 1 + tan ^2 x ) ( 1 + tan^2 x ) (sec x )

OpenStudy (across):

I had to look it up\[\frac{\sin(x)\sec^{m-1}(x)}{m-1}+\frac{m+2}{m-1}\int\sec^{m-2}(x)dx\]

OpenStudy (anonymous):

this is wat i have so far: \[\int\limits_{}^{}\sec^3 x (\sec^2 dx)\] \[\int\limits_{}^{}\sec^3 x (1+\tan^2 x dx)\] \[\int\limits_{}^{}(\sec^3 x +\sec^3 x \tan^2 x )dx\]

OpenStudy (anonymous):

derive tan x = sec^2x

OpenStudy (anonymous):

so do a u sub with u = tan x , du = sec^2 x dx

OpenStudy (anonymous):

\[\int\limits_{}^{}(\sec^3 x) + \int\limits_{}^{}(\sec^3 x \tan^2 x) dx\] can i use u sub after this step

OpenStudy (anonymous):

you will prolly end up with a ( 1 + u ) ^(to some power ) then it a simple integration

OpenStudy (anonymous):

can u show me the steps after the one i just did

OpenStudy (anonymous):

or wat to do?

OpenStudy (anonymous):

hang on...

OpenStudy (anonymous):

im a little rusty

OpenStudy (anonymous):

ughh i hate problems liek this!! XD...requires too much work

OpenStudy (anonymous):

yea, it may have to be broken up into parts I cant get rid of one of the sec^x check out: http://www.wolframalpha.com/input/?i=integrate+sec^5+x

OpenStudy (anonymous):

this IS a reduction formula for powers of sec.

OpenStudy (anonymous):

i did n idont understand it b.c it uses smthng i didnt leard which is teh reduction formula thing

OpenStudy (anonymous):

then again do u mind tellin me wat that is

OpenStudy (anonymous):

or how to use it

OpenStudy (anonymous):

yea, your instructor may not accept a reduction formula answer, it the quick way, but there is a trigonometric solution for this.

OpenStudy (anonymous):

cool i will prob understand that if u explain

OpenStudy (anonymous):

the trig solution i mean

OpenStudy (anonymous):

there is an identity that will lead to an integral of u. anywa here is a like with the reduction formula for secant, and others http://www.sosmath.com/calculus/integration/moretrigpower/moretrigpower.html

OpenStudy (anonymous):

as a matter of fact I bet your textbook has a table of integrals for integrals of powers of secant

OpenStudy (anonymous):

what is the dffnce between sex^n (x) and sec^n (ax)

OpenStudy (anonymous):

a=1

OpenStudy (anonymous):

i wud use the first formula right? sec^5 (x) = sec ^5-2 (x) sec^2(x) = sec^3 (tan^2 x * sec^2 x -1)

OpenStudy (anonymous):

where do u go on from there

OpenStudy (anonymous):

omg i havent done definite integrals!

OpenStudy (anonymous):

no, I mistyped

OpenStudy (anonymous):

∫sec^5dx=1/4sec^3xtanx+3/4∫sec^3xdx

OpenStudy (anonymous):

no definite integrals here

OpenStudy (anonymous):

hmm how did u get 1/4 sec ...etc

OpenStudy (anonymous):

then apply the reduction formula ( or integrate ) \[\int\limits \sec ^ 3 x dx\]

OpenStudy (anonymous):

im confused how u got the previous step thou

OpenStudy (anonymous):

its from the reduction formula, it ends with an integral of ∫sec^3xdx

OpenStudy (anonymous):

i mean teh 1/4 sec part im confused about

OpenStudy (anonymous):

in the reduction formula it 1/(a(n-1) n=5, the power of secant a = 1

OpenStudy (anonymous):

ok so now how did u get 3/4?

OpenStudy (anonymous):

from the reduction formula: (n-2)/(n-1) again n = 5, the power of the secant

OpenStudy (anonymous):

you instructor will be either impressed or put off that you are using the integral tables. :)

OpenStudy (anonymous):

lol...he will b impressed hopefully...i have a test 2mm n im trying to udnerstand probs idk

OpenStudy (anonymous):

yea the trigonometric integrals take practice( obviously )

OpenStudy (anonymous):

srry this is new to me so after taht step wat do u do..u still have a integral of sec^3 dx to deal with

OpenStudy (anonymous):

right! so integrate the sec^x or use the formula again, but this time with n = 3, see?

OpenStudy (anonymous):

sec^x is short, but it indefinite integral has many parts.

OpenStudy (anonymous):

so after using n=3 do u get 1/2 sec x dx

OpenStudy (anonymous):

right ( n-2)/(n-1) this time the formula will end with \[\int\limits \sec x dx\]

OpenStudy (anonymous):

wait but isnt is half secant

OpenStudy (anonymous):

so how is teh integral jsut sec

OpenStudy (anonymous):

no the formula will end ith ∫secxdx, integrate that and you are done.

OpenStudy (anonymous):

\[\int\limits \sec x dx = \ln ( \sec x +tanx ) + C\]

OpenStudy (anonymous):

so teh final answer is: 1/4 sec^3 x tanx + 3/4 ...idk teh rest

OpenStudy (anonymous):

1/4 sec^3 x tanx + 3/4∫sec^3xdx then apply the formula again for ∫sec^3xdx that answer will end with ∫sec xdx integrate ∫secxdx, done!

OpenStudy (anonymous):

so the final asnwer is: 1/4 sec^3 x tanx + 3/8 sec x tanx + tn (secx+tanx) + C

OpenStudy (anonymous):

SRY TAHTS SUPPOOSED TO be ln not tn

OpenStudy (anonymous):

right, distribute the 3/4, I didnt work it all out but it looks right. Those tables are in you textbook right, in the appendix?

OpenStudy (anonymous):

let me check now

OpenStudy (anonymous):

"table of integrals"

OpenStudy (anonymous):

under wat heading shud i looj for te hreduction formulas

OpenStudy (anonymous):

i see basic forms, forms invlving sqrt of (a^2+u^2, forms involving (a^2 - u^2), forms involvung sqrt of (u^2 - a^2), forms involving a+bu, trig forms, inverse trig forms, exp and log forms, hyperbolic forms

OpenStudy (anonymous):

Tables of integrals, Lists of Integrals usually in the appendix, could be on the inside covers.

OpenStudy (anonymous):

yeah i see them in the appenxix...jsut under wat heading of integrals will i fimd teh reduction formulas

OpenStudy (anonymous):

or wher u jsut letting me know taht tehre was a table of integrtals

OpenStudy (anonymous):

that really what we were using, an integration formula, so you can integrate by parts or look up the formula.

OpenStudy (anonymous):

there might be one for ∫sec^n (ax)dx

OpenStudy (anonymous):

might not http://integral-table.com/

OpenStudy (anonymous):

ths is wat is in my bk...sec^n du = 1/n tan u sec^n-2 + (n-2/n-1) *integral of sec^n-2 u du

OpenStudy (anonymous):

i guess tahts anoterh form of te one u showed me

OpenStudy (anonymous):

right that one dont account for a coefficient on x. so, yea, you seem to have it. Im out. ( gotta brush up on my integrals ;) )

OpenStudy (anonymous):

THANK U SOOOOO MUCH!!! PLUS THE WEBSITE LOOKS SUPER HELPFUL =)

OpenStudy (anonymous):

peace

OpenStudy (anonymous):

ii imma go study sum more

OpenStudy (anonymous):

gnite